OpenStudy (anonymous):

A snow cone stand sells crushed, flavored ice in cone-shaped paper cups. The opening of the cup is 8 centimeters wide and the height of the cup is twelve centimeters. What is the approximate volume of the cup itself? two hundred point nine six cubic centimeters six hundred two point eight eight cubic centimeters eight hundred three point eight four cubic centimeters two thousand, four hundred eleven point five two cubic centimeters

OpenStudy (dmb1215):

b

OpenStudy (anonymous):

volume of a cone= pi*r^2*h/3 where r=4cm and h=12

OpenStudy (dmb1215):

4*4 16 *12=192

OpenStudy (dmb1215):

sorry a

OpenStudy (anonymous):

So it's A?

OpenStudy (dmb1215):

yep

OpenStudy (anonymous):

yes it is.